# Mixture of Routers

### v1

@@ 1 Introduction
  • $\bullet$
    We propose a variant of MoR called Consistent Routing Weighting (CRW) to address the impact of sharp, erratic changes in transfer learning, effectively enhancing the model’s stability and generalization ability.
@@ 3.2 Architecture
Table 1: Direct Fine-Tuning Performance comparison of different models on various tasks. We report the accuracy of all the tasks. Higher is better for all metrics. We use the same hyperparameters, which are specified in section LABEL:Hyperparameters. The best results are denoted in bold.
Models SciQA ComQA OpenQA MRPC CoLA RTE Avg.
LoRA 91.01 74.61 76.6 84.41 86.95 84.48 83.01
MoLA 92.94 77.97 78.7 84.52 86.64 86.48 84.54
MoLA + MoR 92.09 77.56 80.0 83.94 85.43 88.09 84.52
MoLA + CRW 93.21 78.79 81.6 85.39 86.77 88.45 85.70
Table 2: Instructional Fine-Tuning Performance comparison of different models on various tasks. CRW is a variant of MoR.
3.3 Consistent Routing Weighting
In transfer learning, models face the challenge of distributional differences between the source and target domains. During the process of using MoR for transfer learning, the model heavily relies on the main router, making it difficult to adjust the distribution of expert routes to adapt to new tasks in the early stages of transfer learning. To address these issues, we propose a variant of MoR, called Consistent Routing Weighting (CRW) , to enhance the model’s stability and generalization ability. This method ensures parameter stability by applying equal weights across different routes, thereby minimizing the risk of overfitting to specific tasks. By preventing the model from overly relying on features from a particular route, we maintain balanced feature representation, which is crucial when handling tasks from different domains. CRW helps smooth the optimization landscape, making the training process more stable and reducing the likelihood of sharp fluctuations in the loss function. Additionally, CRW enhances cross-domain robustness by ensuring a more uniform contribution from each route, making the model less sensitive to domain-specific variations and better equipped to transfer knowledge effectively. This process can be represented as:
$F_{i}(x)^{CRW}=\sum_{i=1}^{N}\frac{1}{N}\cdot r^{i}.$ (7)
@@ 4.2 Main Results
Instructional Fine-Tuning. We first tune LLAMA-2-7B on the instructional tuning dataset using each PEFT method. Then, we fine-tune for all downstream tasks. Instructional tuning effectively evaluates the transfer learning capabilities of each PEFT method. We only compare methods based on LoRA, as they exhibit stronger transfer learning capabilities compared to methods based on prompt tuning. The results, as shown in Table 2, indicate that MoLA+MoR achieved similar results to MoLA on the Instructional Fine-Tuning task, especially on the ScienceQA task where it decreased by 0.85%. However, after employing the MoR variant CRW, there was a noticeable performance improvement, particularly on the CommonsenseQA and OpenbookQA tasks, where MoLA+CRW improved by 2.9% and 0.82% respectively compared to MoLA. On average, there was a 1.16% improvement. Experimental results show that as a plug-in for MoE-style models, MoR significantly boosts the performance of direct fine-tuning, while in the Instructional Fine-Tuning task, the MoR variant CRW performs better in terms of transfer learning.
@@ 4.3.2 Expert Allocation Analysis
Figure 5: Comparison of training loss between MoR and CRW (a variant of MoR) on the OpenBookQA transfer learning task.
4.3.3 Why MoR Performs Poorly on Instruction Fine-tuning Tasks?
MoR dynamically allocates weights via a main router, aiming to flexibly adjust the contribution of each sub-router based on varying task requirements. However, in transfer learning, experimental results show that MoR performs worse than simply using CRW, which contrasts sharply with its superior performance in direct supervised fine-tuning. We recorded the changes in training loss during the early stages of transfer learning, as shown in Figure 5. The results indicate that in the initial phase of transfer learning, CRW exhibits a faster and lower loss reduction. This suggests that equally distributed weights can more efficiently leverage the initial feature representations of multiple routers, allowing for quicker adaptation to the target domain task. In contrast, due to the involvement of the main router, MoR may suffer from suboptimal weight allocation at the beginning, leading to reduced optimization efficiency and slower model convergence. Further analysis reveals that the core reason for this discrepancy lies in the unique characteristics of transfer learning. First, transfer learning faces the challenge of distributional differences between the source and target domains. The main router needs to learn how to allocate weights from limited target domain data. Due to the scarcity of target domain data, the main router is prone to overfitting or developing biases, resulting in suboptimal weight allocation strategies. In comparison, equal weight distribution does not rely on target domain data, avoiding this issue and thus demonstrating stronger robustness. Second, the inherent uncertainty in transfer learning makes it difficult for the main router to accurately capture the importance of each sub-router. If the main router fails to fully understand the characteristics of the target domain, it might mistakenly suppress the contributions of certain critical sub-routers. On the other hand, equal weight distribution ensures equal participation from all sub-routers, fully leveraging the diversity advantage of multi-router architectures.

### v2

@@ 3.2 Architecture
Table 1: Direct Fine-Tuning Performance comparison of different models on various tasks. We report the accuracy of all the tasks. Higher is better for all metrics. We use the same hyperparameters, which are specified in section C. The best results are denoted in bold.
@@ 5 Conclusion
Appendix A Datasets
A.1 NLP Datasets
For evaluation, we adopt the GLUE (General Language Understanding Evaluation) benchmark, which is a widely used collection of datasets designed to assess natural language understanding (NLU) capabilities of models. Specifically, the GLUE benchmark includes several datasets such as COLA Wang et al. (2018), where the task is to determine whether a given sentence is grammatically acceptable; MRPC Dolan and Brockett (2005), which focuses on identifying whether two sentences are semantically equivalent; and RTE Wang et al. (2018), where the goal is to predict if a premise sentence entails a hypothesis sentence. We present the dataset statistics of GLUE in the following table 3.
Dataset Metric #Train #Valid #Test #Label
COLA Mcc 8.5k 1,043 1,063 2
MRPC Acc 3.7k 408 1.7k 2
RTE Acc 2.5k 277 3k 2
Table 3: Dataset Sizes and Evaluation Metrics in the GLUE Benchmark. "Mcc" and "Acc" denote the Matthews correlation coefficient and accuracy
A.2 CR Datasets
Specifically, for the CR tasks, we evaluate ScienceQA Lu et al. (2024), a dataset that tests multi-modal reasoning across text, images, and scientific concepts; CommonsenseQA Talmor et al. (2019), which focuses on answering questions that require everyday commonsense knowledge and reasoning about implicit relationships; and OpenbookQA Mihaylov et al. (2018), designed to evaluate reasoning over explicit scientific facts and principles provided in an open book format. For more details and specific statistics, please refer to Table 4.
Dataset Metric #Train #Valid #Test #Label
SciQA Acc 6,508 - 2,224 4
ComQA Acc 9,740 1,221 - 5
OpenQA Acc 4,957 500 500 4
Table 4: Dataset Sizes and Evaluation Metrics for Science Question Answering Tasks. "Acc" denotes accuracy.
Appendix B Redundancy and Fault Tolerance Theory
Redundancy and Fault Tolerance Theory is a fundamental concept in system design, particularly in engineering, computer science, and reliability analysis. It focuses on enhancing system robustness by incorporating redundant components or mechanisms to ensure continued operation despite failures. The theory is based on the principle that adding extra resources, such as backup systems or duplicate components, can mitigate the impact of faults, thereby improving overall reliability and availability. A key metric in this theory is the system reliability, often modeled using probability:
$R_{sys}=1-(1-R)^{n}$ (7)
where R represents the reliability of an individual component, and n is the number of redundant components. This formula illustrates how redundancy increases system reliability exponentially with additional components. Fault tolerance extends beyond hardware to software systems, employing techniques like error detection, correction codes, and failover mechanisms.
Appendix C Hyperparameters
To facilitate the reproducibility of our experimental results, we have made the hyperparameters used in our experiments publicly available. This includes detailed configurations such as learning rates, batch sizes, optimization algorithms, weight decay, and other relevant settings. By providing this information, we aim to ensure that our experiments can be replicated and validated by other researchers, fostering transparency and enabling further advancements in the field.
         Hyperparameters          Value
         batch_size          64
         micro_batch_size          4
         learning_rate          3e-4
         cutoff_len          256
         lora_r          8
         lora_dropout          0.05
         lora_alpha          16
Table 5: Hyperparameters
Figure 5: Visualization of MoR expert activations, with the total number of experts being 2, 4, 6, and 8 from top to bottom.
Appendix D Visualization of expert allocation
In this section, we present the visualization of expert allocation and activation within the MoR framework, as shown in Figure 5. The similarity in color across each row indicates that the MoR plugin effectively mitigates the issue of improper and uneven expert allocation commonly observed in MoE-style models.